# A Unified Approach to Enforce Non-Negativity Constraint in Neural Network Approximation for Optimal Voltage Regulation

### v2

@@ 2.2 Theoretical Guarantee of Input Convex Neural Networks for Voltage Regulation
There are two requirements in the ICNN, which are special activation functions and non-negativity constraints. Commonly used activation functions, such as ReLU and its variants like leaky ReLU, satisfy both convexity and monotonicity, ensuring nonlinearity for universal NN approximation. However, constraining weights to the positive range limits training flexibility and representation efficiency. For example, [26] uses a processing mechanism following standard NN training, replacing negative weights iteratively, which can cause oscillations in gradients and weights. Alternatively, others embed a gating function into the NN architecture to constrain weights, such as the “clamp function” in PyTorch. This function clamps all elements in input $x$ into the range $[x_{\min},x_{\max}]$. Both strategies are separated from the back-propagation in training, bringing challenges to the convergence of the ICNN. To avoid oscillations in training, we propose implementing the constraint in parallel with training, providing a satisfactory guarantee of a hard constraint for input convexity.
@@ 3.1 Inefficiency of the Duplication Trick in ICNN
Due to the non-negativity constraint on the weights, the ICNN loses significant representation power despite the linear mapping of the “passthrough” layers ${\bm{U}}_{0:k-1}$ being designed to mitigate this issue. [26] uses a duplicate of ${\bm{x}}$ to improve the representation power of the basic ICNN. The negative weights of ${\bm{U}}_{0:k-1}$ in [26] are set to zero, and their negations are set as the weights for corresponding $-{\bm{x}}$. This way, the forward calculation in the new structure has the same result as the calculation in the original ICNN structure. For instance, consider an input pair $[x_{1},x_{2}]^{T}$ and the corresponding weights pair $[w_{1},w_{2}]$, where $w_{2}<0$. Following this method, the new weights pair becomes $[w_{1},0,0,-w_{2}]^{T}$. With the combination of the original inputs and mirroring inputs $[x_{1},x_{2},-x_{1},-x_{2}]$, the result is $(w_{1}x_{1}+w_{2}x_{2})$, which equals the inner product of $[w_{1},w_{2}]$ and $[x_{1},x_{2}]$.
@@
From the perspective of weights updating, the network will reset all weights of $-{\bm{x}}$ after each iteration, so these weights ${\bm{U}}_{0:k-1}^{(-)}$ are only considered in the forward calculation, not in the back-propagation. Moreover, since all the weights of ${\bm{x}}$, i.e., ${\bm{U}}_{0:k-1}^{(+)}$, are non-negative after each iteration, the negation of their negative value, which is the weights in ${\bm{U}}_{0:k-1}^{(-)}$ will be small in the new iteration and to the final iteration.

### v3

@@ 2.2 Theoretical Guarantee of Input Convex Neural Networks for Voltage Regulation
There are two requirements in the ICNN, which are special activation functions and non-negativity constraints. Commonly used activation functions, such as ReLU and its variants like leaky ReLU, satisfy both convexity and monotonicity, ensuring nonlinearity for universal NN approximation. However, constraining weights to the positive range limits training flexibility and representation efficiency. For example, [26] uses a processing mechanism following standard NN training, replacing negative weights iteratively, which can cause oscillations in gradients and weights. Alternatively, others embed a gating function into the NN architecture to constrain weights, such as the “clamp function” in PyTorch. This function clamps all elements in input $x$ into the range $x_{\min},x_{\max}]$. Both strategies are separated from the back-propagation in training, bringing challenges to the convergence of the ICNN. To avoid oscillations in training, we propose implementing the constraint in parallel with training, providing a satisfactory guarantee of a hard constraint for input convexity.
@@ 3.1 Inefficiency of the Duplication Trick in ICNN
Due to the non-negativity constraint on the weights, the ICNN loses significant representation power despite the linear mapping of the “passthrough” layers ${\bm{U}}_{0:k-1}$ being designed to mitigate this issue. [26] uses a duplicate of $\bm{x}$} to improve the representation power of the basic ICNN. The negative weights of ${\bm{U}}_{0:k-1}$ in [26] are set to zero, and their negations are set as the weights for corresponding $\bm{x}$}. This way, the forward calculation in the new structure has the same result as the calculation in the original ICNN structure. For instance, consider an input pair $x_{1},x_{2}]^{T}$ and the corresponding weights pair $w_{1},w_{2}]$, where $w_{2}<0$. Following this method, the new weights pair becomes $w_{1},0,0,-w_{2}]^{T}$. With the combination of the original inputs and mirroring inputs $x_{1},x_{2},-x_{1},-x_{2}]$, the result is $(w_{1}x_{1}+w_{2}x_{2})$, which equals the inner product of $w_{1},w_{2}]$ and $x_{1},x_{2}]$.
@@
From the perspective of weights updating, the network will reset all weights of $\bm{x}$} after each iteration, so these weights ${\bm{U}}_{0:k-1}^{(-)}$ are only considered in the forward calculation, not in the back-propagation. Moreover, since all the weights of $\bm{x}$}, i.e., ${\bm{U}}_{0:k-1}^{(+)}$, are non-negative after each iteration, the negation of their negative value, which is the weights in ${\bm{U}}_{0:k-1}^{(-)}$ will be small in the new iteration and to the final iteration.